# Unification of Exceptional Points and Transmission Peak Degeneracies in a Highly Tunable Magnon-Photon Dimer

### v2

@@ III.0.1 Eigenspectrum Analysis
$\lambda_{0}$ $=\left(\frac{\Delta_{\kappa}}{2}-\frac{\kappa_{c}}{2}\right)+\mathrm{i}\left(\frac{\Delta_{f}}{2}-f_{c}+f_{d}\right),$ (6)
$\Delta_{\lambda}$ $=\sqrt{-\Delta_{f}^{2}+2\mathrm{i}\Delta_{f}\Delta_{\kappa}+\Delta_{\kappa}^{2}-4J^{2}e^{\mathrm{i}\phi}},$ (7)
$\lambda_{\pm}$ $=\lambda_{0}\pm\frac{\Delta_{\lambda}}{2}.$ (8)
The quantity $\Delta_{\lambda}$ defines the complex spectral splitting between the eigenvalues, and, remarkably, appears throughout the analysis for EPs, Petermann noise factor, and even TPDs in the form $\Delta_{\lambda}^{2}$. Since the radicand of Eq. 7 is generally complex, the square root is computed using the principal value, which we can define in polar form as $\Delta_{\lambda}=\sqrt{|\Delta_{\lambda}|^{2}}\,e^{\mathrm{i}\arg(\Delta_{\lambda}^{2})/2}$, with the argument $\arg(\Delta_{\lambda}^{2})\in(-\pi,\pi]$. This choice ensures continuity of $\Delta_{\lambda}$ in parameter space except at branch points along the zero contour $\Delta_{\lambda}=0$. This zero contour is a curve along which the two eigenvalues are degenerate in real and imaginary components. To determine whether these degeneracies are ordinary crossings or EPs, we examine the eigenvector coalescence, quantified by the Petermann noise factor.
The Petermann noise factor is defined as $K_{j^{\prime}j}=\langle L_{j^{\prime}}|L_{j}\rangle\langle R_{j}|R_{j^{\prime}}\rangle$, where $\langle L_{j}|$ and $|R_{j}\rangle$ denote the left and right eigenvectors of $\bm{A}$ [28, 7]. For convenience, we use the mean diagonal Petermann noise factor $\bar{K}_{2}=\frac{1}{2}(K_{11}+K_{22})$, whose deviation from unity measures eigenvector non-orthogonality and diverges at EPs [58]. While originally formulated to describe excess spontaneous emission in unstable lasers [28], the Petermann noise factor now serves more generally as a predictor of noise enhancement in non-Hermitian systems [27, 59]. For $\bm{A}$ in Eq. 5, the mean diagonal Petermann noise factor is
$\bar{K_{2}}=\frac{\Delta_{f}^{2}+\Delta_{\kappa}^{2}+|\Delta_{\lambda}|^{2}+4J^{2}}{2|\Delta_{\lambda}|^{2}}.$ (9)
When $\Delta_{f}=\Delta_{\kappa}=0$, Eq. 9 evaluates to unity for all values of $\phi$ and $J$, yielding purely orthogonal eigenvectors. In this regime, the only possible spectral degeneracies are diabolic points (DPs), which correspond to real-eigenvalue crossings and require $J=0$ to occur. In contrast, introducing nonzero $\Delta_{f}$ and $\Delta_{\kappa}$ enables eigenvector non-orthogonality and thus eigenvector coalescence. Equation 9 diverges as $|\Delta_{\lambda}|\to 0$, while the numerator remains finite, signaling the presence of an EP. For a given $\phi\in[0,2\pi)$, solving $|\Delta_{\lambda}|=0$, equivalent to finding where $\text{Re}(\Delta_{\lambda})=\text{Im}(\Delta_{\lambda})=0$, yields a continuous manifold of EPs in parameter space, called an exceptional surface [60]. This exceptional surface is parameterized by
$\Delta_{\kappa}=\pm 2J\cos\left(\frac{\phi}{2}\right),\quad\Delta_{f}=\pm 2J\sin\left(\frac{\phi}{2}\right).$ (10)
For our experiments, we analyze two distinct EP configurations along the exceptional surface, namely the reciprocal phase EP (RP-EP) and anti-reciprocal phase EP (ARP-EP), given by
ARP-EP: $\phi=\pi,\Delta_{f}^{\text{EP}}=\pm 2J,\Delta_{\kappa}=0,$
RP-EP: $\phi=0,\Delta_{f}=0,\Delta_{\kappa}^{\text{EP}}=\pm 2J.$
Moreover, the Supplementary Information explores a hybrid EP at $\phi=\pi/2$, where the coupling is truly non-reciprocal.
Figure 2: Geometry of EPs and TPDs. a, d Absolute imaginary eigenvalue splitting $|\text{Im}(\Delta_{\lambda})|$, with cyan curves representing the zero contour of the discriminant $\text{Disc}=0$, separating the parameter space into single peak (between cyan contours horizontally) and two peak (outside cyan contours horizontally) regions. The magenta curves are the contour $q=0$, representing the path an experiment must take to cross a TPD optimally, equivalent to $\Delta_{f}=0$ ($\Delta_{\kappa}=0$) for the RP-TPD (ARP-TPD). The TPDs (red circles) occur when the magenta and cyan curves intersect, and do not coincide with the EPs (red crosses) for these choices of $\kappa_{c}/J$. b, e depict the real eigenvalue splitting $\text{Re}(\Delta_{\lambda})$, with lime curves representing the contour $\text{Re}(\Delta_{\lambda})=\bar{\kappa}$, signaling an instability transition. The unstable region, where the linear model is insufficient, is depicted in a gray colorscale. c, f depict the mean of the diagonal Petermann noise factor $\bar{K}_{2}$, which diverges at the EPs, but not at the TPDs. Both control parameter sets match the operating points used in Fig. 3, with $\kappa_{c}/J\approx 0.69$ for panels a–c and $\kappa_{c}/J\approx 0.092$ (d–f). Written in terms of dimensionless parameters, $\Delta_{f}/J,\Delta_{\kappa}/J,\bar{\kappa}/J$, the shape of the contours becomes independent of the value of $J$ (see Supplementary Information).
III.0.2 Transmission Spectrum Analysis
Equation 2 governs the time dynamics of the driven two-mode system, with a general solution given by $\bm{\alpha}(t)=e^{\bm{A}t}\bm{\alpha}(0)+\bm{A}^{-1}(e^{\bm{A}t}-\bm{I})\bm{B}u$ for time-independent drive $u$ and nonsingular $\bm{A}$. During experiments, we perform a homodyne readout of the transmission signal, which captures the steady-state mode amplitudes after transient dynamics have ceased. To model this behavior, we set $\dot{\bm{\alpha}}=0$ in Eq. 2 and solve for the steady-state response $\bm{\alpha}_{\mathrm{ss}}$ with
$\bm{\alpha}_{\mathrm{ss}}$ $=$ $-\bm{A}^{-1}\bm{B}\,u,$ (11)
${\beta}_{\mathrm{ss}}$ $=$ $|\bm{C}\bm{\alpha}_{\mathrm{ss}}|^{2},$ (12)
where $\beta_{\mathrm{ss}}$ is proportional to the detected power. Equation 11 is only valid when the system is stable, given by the Hurwitz stability criterion, which requires all real parts of the eigenvalues $\bm{A}$ to be negative. When $\bm{A}$ becomes unstable, the dynamics are eventually stabilized by saturation of amplifiers, leading to nonlinear phenomena, characterized in previous work but not considered here [55] (See Supplementary Information for details). Transmission peak frequencies correspond to the critical points of $\beta_{ss}$, obtained by solving $\frac{\partial}{\partial f_{d}}(\beta_{\mathrm{ss}})=0$ for $f_{d}$. This condition yields a nonlinear equation for $f_{d}$, which can be expressed as a depressed cubic in terms of an auxiliary variable $x$ of the form
00 0 $=x^{3}+px+q,$ (13)
$p$ $=\frac{\bar{\kappa}^{2}+\text{Re}(\Delta_{\lambda}^{2})}{4},$ (14)
$q$ $=\frac{\bar{\kappa}\text{Im}(\Delta_{\lambda}^{2})}{8}.$ (15)
The three roots of Eq. 13 are the critical points of $\beta_{ss}$ and, remarkably, are related to the eigenspectrum of $\bm{A}$ in Eq. 5 through $\Delta_{\lambda}^{2}$ in Eq. 7.
When Eq. 13 admits three real roots, the outermost roots (the largest and smallest), denoted $\nu_{\pm}^{\text{Root}}$, are used to find the transmission peak locations, while the intermediate root, denoted $\eta^{\text{Root}}$, is used to find the local minimum between the two peaks. When there is no splitting, Eq. 13 admits one real root, denoted $\nu_{0}^{\text{Root}}$, corresponding to the single peak frequency. Solving Eq. 13 for a closed form solution of the roots in terms of $p,q$ can either be done using Cardano’s method, which involves square and cube roots, or through a trigonometric solution first posed by François Viète in 1615 [61, 62, 63], which yields a solution for $\nu_{\pm}^{\text{Root}}$ as
$\nu_{\pm}^{\text{Root}}$ $=2\sqrt{\frac{-p}{3}}\,\cos\!\left(\theta\pm\frac{\pi}{3}\right),$ (16)
$\theta$ $=\frac{\pi}{3}+\frac{1}{3}\arccos\!\left(\frac{3q}{2p}\sqrt{\frac{-3}{p}}\right),$ (17)
and a solution for $\eta^{\text{Root}}$ given by
$\eta^{\text{Root}}$ $=2\sqrt{\frac{-p}{3}}\,\cos\!\left(\theta\right).$ (18)
The expression for the splitting between the two peaks, $\Delta_{\nu}^{\text{Root}}=\nu_{+}^{\text{Root}}-\nu_{-}^{\text{Root}}$, admits a compact expression in terms of $\theta$ in Eq. 17 as
$\Delta_{\nu}^{\text{Root}}$ $=-2\sqrt{-p}\,\sin\left(\theta\right).$ (19)
Equations 16 and 19 are undefined at $p=0$, and are otherwise only valid in regions where the discriminant, expressed algebraically as
$\text{Disc}\equiv_{\text{algebraic}}-4p^{3}-27q^{2},$ (20)
is positive, signifying three real root solutions to Eq. 13. In contrast, when $\text{Disc}<0$, Eq. 16 and 19 are invalid, as there is only one real root, $\nu_{0}^{\text{Root}}$, given by
$\nu_{0}^{\text{Root}}=\begin{cases}-2\dfrac{|q|}{q}\sqrt{\dfrac{-p}{3}}\cosh\!\left[\dfrac{1}{3}\,\mathrm{arcosh}\left(\dfrac{-3|q|}{2p}\sqrt{\dfrac{-3}{p}}\right)\right],&\text{if }p<0,\\[5.81247pt] -2\sqrt{\dfrac{p}{3}}\sinh\!\left[\dfrac{1}{3}\,\mathrm{arsinh}\left(\dfrac{3q}{2p}\sqrt{\dfrac{3}{p}}\right)\right],&\text{if }p>0.\end{cases}$ (21)
Naturally, the condition $\text{Disc}=0$ marks the boundary between the single-peak and split-peak regimes. This boundary can be represented as a zero-contour in $\Delta_{f}$-$\Delta_{\kappa}$ space, plotted in Fig. 2a, d. However, $\text{Disc}=0$ alone does not define the locations of TPDs, as it is not guaranteed that $\Delta_{\nu}^{\text{Root}}\rightarrow 0$ at $\text{Disc}=0$. This can be illuminated through the geometric representation of the discriminant
$\text{Disc}\equiv_{\text{geometric}}(r_{1}-r_{2})^{2}(r_{2}-r_{3})^{2}(r_{3}-r_{1})^{2},$ (22)
where $r_{1},r_{2},r_{3}$ are the roots of Eq. 13. For $\text{Disc}<0$, one of the roots is real, given by $\nu_{0}^{\text{Root}}$, and the other two roots are a complex conjugate pair, not considered here. For $\text{Disc}>0$, $r_{1},r_{2},r_{3}$ are $\nu_{-}^{\text{Root}},\eta^{\text{Root}},\nu_{+}^{\text{Root}}$ respectively. Finally, at $\text{Disc}=0$, a multi-root degeneracy occurs, corresponding to the coalescence of either $\nu_{+}^{\text{Root}}$ or $\nu_{-}^{\text{Root}}$ with the central root $\eta^{\text{Root}}$, but not necessarily all three required for a TPD.
To differentiate between $\text{Disc}=0$ causing a double-root degeneracy or a true TPD, an additional condition is required. In Eq. 20, enforcing $\text{Disc}=0$ and $q=0$ causes $p=0$ as well. Then, $p=q=0$ simplifies Eq. 13 to $x^{3}=0$, a triple degeneracy. Thus, TPDs occur precisely at intersections of the zero contours defined by $\text{Disc}=0$ and $q=0$, which are plotted in Fig. 2. The zero contour $q=0$ is the sensor optimal path and is plotted in Fig. 2a,c,d,f. The optimal path ($q=0$) is the parameterized curve treated as the independent variable during experimental operation, and is $\Delta_{f}=0$ ($\Delta_{\kappa}=0$) for the RP (ARP) configuration. In the Supplementary Information, we comment on how this relates to choosing a TPD based on the type of sensor being built and the added complexity of $\phi\not\in\{0,\pi\}$ TPDs, where the optimal paths become curved contours requiring precise, simultaneous control of $\Delta_{f}$ and $\Delta_{\kappa}$.
III.0.3 Unifying EPs, TPDs, and Instability
To make a concrete connection between the theoretical treatment in the previous section and our experimental results, we begin by converting the roots, $\nu_{\pm}^{\text{Root}}$ to frequencies observed experimentally using
$\nu_{\pm,0}=\nu_{\pm,0}^{\text{Root}}+f_{c}-\frac{\Delta f}{2},$ (23)
where we can see, naturally, that $\Delta_{\nu}^{\text{Root}}=\Delta_{\nu}$.
To unify EPs and TPDs, we first consider the cases where TPDs appear exactly at EPs. For the RP configuration, the condition $\kappa_{c}=2J,\kappa_{y}=-2J$ forms a TPD-EP pair with balanced loss and gain, with $\nu_{\pm}=|\text{Im}(\lambda_{\pm})|$ locally around the TPD. For the ARP configuration, the condition $\kappa_{c}=0,\kappa_{y}=0$ forms TPD-EP pairs at $\Delta_{f}=\pm 2J$, with $\nu_{\pm}=|\text{Im}(\lambda_{\pm})|$ globally. A geometric interpretation of the correspondence between peak locations and imaginary eigenvalues is presented in the Supplementary Information. Next, we generalize the correspondence between TPDs, $\nu_{\pm}$, EPs, and $|\text{Im}(\lambda_{\pm})|$, to the cases where both modes are dissipative ($\kappa_{c,y}>0$). In this more typical scenario, the peak locations are distinct from the imaginary eigenvalues, and the TPDs shift in parameter space away from the EPs with $\kappa_{c}/J$.
Figure 3: Experimentally Configurable EPs and TPDs. a,b Representative RP- and ARP-TPD configurations with different $\kappa_{c}/J$ ratios. i Imaginary components of $\lambda_{\pm}$, along with $\nu_{\pm}$. ii Real components of $\lambda_{\pm}$, with $\max(\lambda_{\pm})=0$ signifying an instability transition (with a horizontal dashed line indicating 0). iii, Experimentally measured transmission peak locations (error bars represent propagation of uncertainty from $\sigma_{\text{fit}}(f_{c,y}),\sigma_{\text{fit}}(\kappa_{c,y})$, see Methods) along with predicted $\Delta_{f,\kappa}^{\text{EP}}$, $\Delta_{f,\kappa}^{\text{TPD}}$, and $\Delta_{f}^{\text{NL}}$. Uncertainties $\sigma_{\text{exp}}(\Delta_{f,\kappa}^{\text{EP}})$, $\sigma_{\text{exp}}(\Delta_{f,\kappa}^{\text{TPD}})$, $\sigma_{\text{exp}}(\Delta_{f}^{\text{NL}})$, are calculated using $\sigma_{\text{exp}}(J)$ and $\sigma_{\text{exp}}(\kappa_{c})$ (see Methods). Shaded region represents distribution of possible $\nu_{\pm}$, $\mathcal{D}_{\text{exp}}(\nu_{\pm})$, with standard deviation $\sigma_{\text{exp}}(\nu_{\pm})$, using Monte Carlo uncertainty propagation of $\sigma_{\text{exp}}(J),\sigma_{\text{exp}}(\kappa_{c})$ (see Methods).
The locations of EPs and TPDs, along with contours for splitting and instability thresholds, are visualized for a specific $\kappa_{c}/J$ ratio in Fig. 2a, b, d, e for the RP configuration ($\phi=0,\kappa_{c}/J\approx 0.69$)(Fig.2a, b, c) and the ARP configuration ($\phi=\pi,\kappa_{c}/J\approx 0.092$)(Fig.2d, e, f). In the ARP configuration, the EP occurs at $\Delta_{f}^{\text{EP}}=\pm 2J$, while the TPD occurs at $\Delta_{f}^{\text{TPD}}=\sqrt{4J^{2}+{\kappa}_{c}^{2}}$ (Fig. 2d). The system becomes unstable and nonlinear at $\Delta_{f}^{\text{NL}}=\sqrt{4J^{2}-{\kappa}_{c}^{2}}$, provided $0\leq\kappa_{c}<2J$ (Fig. 2e). The instability transition is determined using the condition $\text{Re}(\Delta_{\lambda})=\bar{\kappa}$. In the RP configuration, the RP-EP lies at $\Delta_{\kappa}^{\text{EP}}=\pm 2J$, but the RP-TPDs occur at $\Delta_{\kappa}^{\text{TPD}}=\frac{\kappa_{c}}{2}\pm\frac{\sqrt{8J^{2}-\kappa_{c}^{2}}}{2}$, provided $0\leq\kappa_{c}<\sqrt{8}J$ (Fig. 2d). The TPD can be displaced from the EP in parameter space, significantly reducing the Petermann noise factor at the RP-TPD (Fig. 2c), compared to the ARP-TPD in Fig. 2f. The instability transition occurs at $\Delta_{\kappa}^{\text{NL}}=\kappa_{c}$ for $\kappa_{c}\leq 2J$, and $\Delta_{\kappa}^{\text{NL}}=\frac{4J^{2}+\kappa_{c}^{2}}{2\kappa_{c}}$ for $\kappa_{c}>2J$. Ultimately, this separation between the EPs and TPDs is crucial for interpreting the experimental data in Fig. 3 and understanding the functional role of EPs versus TPDs in sensor applications.
IV Results
Here, we report experimental data for both RP and ARP configurations, selecting deliberately disparate values of $J$ to highlight the contrasting behavior between TPDs that lie close to or far from their corresponding EPs. To ensure the validity of our linear model, we select TPDs that involve purely dissipative dynamics. For the RP configuration, we select a representative small $J$ value of $J=1.01(4)$ MHz with $\kappa_{c}=0.70(1)$ MHz, yielding a ratio $\kappa_{c}/J=0.69(3)$. As shown in Fig. 2c, this choice of RP-TPD is significantly displaced from the RP-EP. For the ARP configuration, we select a representative large $J$ value of $J=8.15(4)$ MHz, with $\kappa_{c}=0.752(8)$ MHz. This choice of $\kappa_{c}/J=0.092(1)$ places the TPD, EP, and instability transition within a 50 kHz window (Fig. 2f).

### v3

@@ II.1 Eigenspectrum Analysis
$\tilde{\lambda}_{0}$ $=\left(\frac{\tilde{\Delta}_{\kappa}}{2}-\frac{\tilde{\kappa}_{c}}{2}\right)+\mathrm{i}\left(\frac{\tilde{\Delta}_{f}}{2}-\tilde{f}_{c}+\tilde{f}_{d}\right),$ (4)
$\tilde{\Delta}_{\lambda}$ $=\sqrt{-\tilde{\Delta}_{f}^{2}+2\mathrm{i}\tilde{\Delta}_{f}\tilde{\Delta}_{\kappa}+\tilde{\Delta}_{\kappa}^{2}-4e^{\mathrm{i}\phi}},$ (5)
$\tilde{\lambda}_{\pm}$ $=\tilde{\lambda}_{0}\pm\frac{\tilde{\Delta}_{\lambda}}{2}.$ (6)
where $\text{Re}(\tilde{\lambda})$ is related to dissipation or gain, and $\text{Im}(\tilde{\lambda})$ is related to oscillation frequency. This convention is opposite to the typical Hamiltonian-based formulation, but consistent with standard practice in input–output theory [35]. The quantity $\tilde{\Delta}_{\lambda}$ defines the complex spectral splitting between the eigenvalues. Figure 2 i(ii) depicts $|\text{Im}(\tilde{\Delta}_{\lambda})|$ ($\text{Re}(\tilde{\Delta}_{\lambda})$) on the $\tilde{\Delta}_{\kappa}$-$\tilde{\Delta}_{f}$ plane.
EPs are defined by the coalescence of eigenvectors and eigenvalues. The mean diagonal Petermann noise factor, $\bar{K}_{2}=\frac{1}{2}(K_{11}+K_{22})$, is used to quantify eigenvector non-orthogonality. $\bar{K}_{2}=1$ when eigenvectors are orthogonal, and diverges at EPs [44, 45, 10]. For $\bm{\tilde{A}}$ in Eq. 3, we find
$\bar{K}_{2}=\frac{\tilde{\Delta}_{f}^{2}+\tilde{\Delta}_{\kappa}^{2}+|\tilde{\Delta}_{\lambda}|^{2}+4}{2|\tilde{\Delta}_{\lambda}|^{2}}.$ (7)
For a given $\phi\in[0,2\pi)$, solving $|\tilde{\Delta}_{\lambda}|=0$ for $\tilde{\Delta}_{\kappa},\tilde{\Delta}_{f}$ simultaneously yields a continuous manifold of EPs in parameter space, called an exceptional surface [46], parameterized by
$\tilde{\Delta}_{\kappa}=\pm 2\cos\left(\frac{\phi}{2}\right),\quad\tilde{\Delta}_{f}=\pm 2\sin\left(\frac{\phi}{2}\right).$ (8)
At each value of $\phi$, there are two EPs (“primary” and “other”), plotted in Fig. 2, where $\text{Re}(\tilde{\Delta}_{\lambda})=\text{Im}(\tilde{\Delta}_{\lambda})=0$, with “primary” EPs investigated experimentally in Fig. 3. In the vicinity of EPs, $\bar{K}_{2}$ diverges, corresponding to noise enhancement at those parameter regimes [9, 47]. Moving away from the EPs in the $\tilde{\Delta}_{\kappa}$-$\tilde{\Delta}_{f}$ plane reduces $\bar{K}_{2}$, reaching a minimum value of $\bar{K}_{2}=1$ between the two EPs. This line of minimal $\bar{K}_{2}$ is the $\tilde{\Delta}_{f}$ axis for $\phi=0$ (Fig. a.iii-iv), the $\tilde{\Delta}_{\kappa}$ axis for $\phi=\pi$ (Fig. b.iii-iv), and the line $\tilde{\Delta}_{f}=-\cot(\phi/2)\tilde{\Delta}_{\kappa}$ for all other values of $\phi$, with $\tilde{\Delta}_{f}=-\tilde{\Delta}_{\kappa}$ at $\phi=\pi/2$ (Fig. c.iii-iv).
II.2 Transmission Spectrum Analysis
Transmission peak frequencies are found from the critical points of $\tilde{\beta}_{ss}$, obtained by solving $\frac{\partial}{\partial\tilde{f}_{d}}(\tilde{\beta}_{\mathrm{ss}})=0$ for $\tilde{f}_{d}$. This condition yields a nonlinear equation for $\tilde{f}_{d}$, which can be expressed as a depressed cubic in terms of an auxiliary variable $f$ of the form
00 0 $=f^{3}+\tilde{p}f+\tilde{q},$ (9)
$\tilde{p}$ $=\frac{(\tilde{\kappa}_{c}-\tilde{\Delta}_{\kappa})^{2}+\text{Re}(\tilde{\Delta}_{\lambda}^{2})}{4},$ (10)
$\tilde{q}$ $=\frac{(\tilde{\kappa}_{c}-\tilde{\Delta}_{\kappa})\,\text{Im}(\tilde{\Delta}_{\lambda}^{2})}{8}.$ (11)
Equation 9 can either admit a single real root, $\tilde{\nu}_{0}^{\text{Root}}$, or three real roots, $\tilde{\nu}_{\pm}^{\text{Root}},\tilde{\eta}^{\text{Root}}$. All roots can be converted into experimentally relevant extrema locations by
$\tilde{\nu}_{\pm,0}$ $=\tilde{\nu}_{\pm,0}^{\text{Root}}+\tilde{f}_{c}-\frac{\tilde{\Delta}_{f}}{2},$ (12)
$\tilde{\eta}$ $=\tilde{\eta}^{\text{Root}}+\tilde{f}_{c}-\frac{\tilde{\Delta}_{f}}{2}.$ (13)
When Eq. 9 admits a single real root, $\tilde{\nu}_{0}$ is the single peak frequency. Otherwise, $\tilde{\nu}_{\pm}$ are the two peak frequencies, with $\tilde{\eta}$ representing the frequency of the local minima between the two peaks. The splitting between the peaks is $\tilde{\Delta}_{\nu}=\tilde{\nu}_{+}-\tilde{\nu}_{-}$, analogous to $\tilde{\Delta}_{\lambda}$ from Eq. 5. The analysis of the roots of Eq. 9 is extensively studied in the Supplementary Information, including a closed form solution for the peak locations using a trigonometric solution posed by François Viète in 1615 [48, 49, 50].
TPDs are defined by the coalescence of the two peak frequencies. Mathematically, we locate TPDs by solving $\mathrm{Disc}=\tilde{q}=0$ for $\tilde{\Delta}_{f},\tilde{\Delta}_{\kappa}$ simultaneously, where $\mathrm{Disc}$ is the discriminant of Eq. 9. This surface of TPDs has a significantly more complicated solution than the exceptional surface in Eq. 8, and is fully parameterized in the Supplementary Information. The critical takeaway from this analysis is that the surface of TPDs is dependent on both $\tilde{\kappa}_{c}$ and $\phi$. Depending on the choice of $\tilde{\kappa}_{c}$ and $\phi$, there can be two, four, or six TPDs. This dependence on both parameters is further investigated with theory in Fig. 2 and experiment in Fig. 3.
To observe the symmetric square root splitting of eigenvalues and transmission peaks characteristic of EPs and TPDs, a path must be chosen in the $\tilde{\Delta}_{\kappa}$-$\tilde{\Delta}_{f}$ plane, which becomes the independent variable that ensures symmetric square root splitting. This path is defined by the contour $\tilde{q}=0$ and is plotted in Fig. 2iii, iv. Following $\tilde{q}$ through the TPD ensures $\text{Re}(\tilde{\Delta}_{\lambda})=0$ throughout the sweep (Fig. 2ii), such that both peaks split with symmetric amplitudes. For any value of $\phi$, there are always $\tilde{q}=0$ contours that smoothly intersect both an EP and a TPD, which we probe using our experimental apparatus in Fig. 3. For $\phi=0$, this experimental sweep path is the $\tilde{\Delta}_{\kappa}$ axis (Fig. 3a-b), or the $\tilde{\Delta}_{f}$ axis for $\phi=\pi$ (Fig. 3c-d). For intermediate values of $\phi$, the sweep path becomes a hyperbola defined by $\tilde{\Delta}_{\kappa}\tilde{\Delta}_{f}=2\sin(\phi)$ (Fig. 2c.iii-iv), requiring simultaneous control of $\tilde{\Delta}_{f}$ and $\tilde{\Delta}_{\kappa}$ (Fig. 3e-f) (Methods).
When the sweep path $\tilde{q}=0$ intersects an EP, the eigenvalues split with a square root profile, not necessarily observed in transmission. When $\tilde{q}=0$ intersects the zero-discriminant contour $\mathrm{Disc}=0$, a TPD occurs and the transmission peaks split with a square root profile. When $\tilde{q}=0$ intersects the instability contour, defined as $\text{Re}(\tilde{\Delta}_{\lambda})=(\tilde{\kappa}_{c}-\tilde{\Delta}_{\kappa})$ (Fig. 2iii-iv), the system becomes unstable, and the steady state transmission is no longer governed by Eq. 2. Remarkably, there can be up to 4 other TPDs on $\tilde{q}=0$ contours that do not intersect EPs. We refer to these as rogue TPDs, which are depicted in Fig. 2iii-iv.
Figure 3: Surveying a Surface of EPs and TPDs. Representative EP and TPDs for various $\phi$ and $\tilde{\kappa}_{c}$. i depicts imaginary eigenvalue splitting at the EPs, and transmission peak splitting at the TPDs for $\phi=0$ (a-b), $\phi=\pi$ (c-d) and $\phi=\pi/2$ (e-f). The TPD moves with $\tilde{\kappa}_{c}$, with the left (right) column surveying smaller (larger) $\tilde{\kappa}_{c}$ values that place the TPD and EP closer to (further from) each other. ii depicts real eigenvalues, which also split at EPs, and cross 00 at an instability transition. iii Experimentally measured transmission peak locations overlaid atop theory, with error bars representing fit uncertainty propagated from $\sigma(\tilde{\Delta}_{f}),\sigma(\tilde{\Delta}_{\kappa})$ (Methods). Error bars are scaled by $5$ to improve visibility, most error bars are contained within the data points.
III Experimental Results
The locations of EPs and TPDs, denoted $\tilde{\Delta}_{f,\kappa}^{\mathrm{EP}},\tilde{\Delta}_{f,\kappa}^{\mathrm{TPD}}$ with respect to the independent variable set by sweeping $\tilde{q}=0$ are visualized in Fig. 3. We report the theory and results of EPs and TPDs explored at $\phi=0,\pi,\pi/2$, with representative small and large $\tilde{\kappa}_{c}$ to survey the surface of EPs and TPDs. To ensure the validity of our linear model, we select “primary” EPs and TPDs that involve purely dissipative dynamics (Fig. 2).